F2 F4 F9 Accounting and the Business Environment 57 CAsH-I6, 400 AIR - 1,200 800 > Continuing Problen Problem Pi-t i he fist problm n a sqguence chas beginu an couning opcde The 4200 cycle is continued in Chapter 2 and completed in Chapter 5 Pl-54 Using the accounting equation for transaction analyis, prparing naialA700 100 statements, and calculating return on assets (ROA) Davls Consulting began operations and completed the following transactions during December, 2014 400 RE Dec. 2 Stockholders contributed $18,000 cash in exchange for common stock 2 Paid monthly office rent, $550 3 Paid cash for a compute. $1,800. This equipment is expected to remain in service for five years. Purchased office furniture on account, $4,200. The furniture should last for five years. 4 5 Purchased office supplies on account, $900 9 Performed consulting service for a client on account, $1,500 12 Paid utilities expenses, $250 18 Performed service for a client and received cash of $1,100. 21 Received $1,400 in advance for client service to be performed in the Ber e 800 180 future. (This increases the Unearned Revenue account which is a lability. This account will be explained in more detail in Chapter 2.) Hired an administrative assistant to be paid $2,055 on the 20th day of each month. The secretary begins work immediately 21 26 Paid $400 on account. 28 Collected $300 on account. 30 Cash dvidends of $1,400 were paid to stockholders. 1. Analyse the effects of Davis Consultings transactions on the accounting equation. Use the format of Exhibit 1-5, and include these headings: Cash Accounts Receivable; Office Supplies Equipment; Furniture, Accounts Payable; Unearned Revenue; Common Stock; Dividends; Service Revenue Rent Expense, and Uilities Expense. 2. Prepare the income statement of Davis Consulting for the month ended 3. Prepare the statement of retained carnings for the month ended December 31, 4. Prepare the balance sheet as of December 31, 2014. 5aethe rturn on aes forA December 31, 2014. 2014 NT
Answer #1

1.

Date Assets Liabilities Stockholders' Equity
Cash + Accounts Receivable + Office Supplies + Equipment + Furniture = Accounts Payable + Unearned Revenue + Common Stock - Dividends + Service Revenue - Rent Expense - Utilities Expense
Dec. 2 18000 + + + + = + + 18000 - + - -
2 -550 + + + + = + + - + - 550 -
3 -1800 + + + 1800 + = + + - + - -
4 + + + + 4200 = 4200 + + - + - -
5 + + 900 + + = 900 + + - + - -
9 + 1500 + + + = + + - + 1500 - -
12 -250 + + + + = + + - + - - 250
18 1100 + + + + = + + - + 1100 - -
21 1400 + + + + = + 1400 + - + - -
21 + + + + = + + - + - -
26 -400 + + + + = -400 + + - + - -
28 300 + -300 + + + = + + - + - -
30 -1400 + + + + = + + - 1400 + - -
Bal. 16400 + 1200 + 900 + 1800 + 4200 = 4700 + 1400 + 18000 - 1400 + 2600 - 550 - 250

2.

Davis Consulting
Income Statement
For the Month Ended December 31, 2014
Revenue:
Service revenue 2600
Expenses:
Rent expense 550
Utilities expense 250
Total expenses 800
Net income 1800

3.

Davis Consulting
Statement of Retained Earnings
For the Month Ended December 31, 2014
Balance, December 1, 2014 0
Net income for the month 1800
Less: Dividends 1400 400
Balance, December 31, 2014 400

4.

Davis Consulting
Balance Sheet
December 31, 2014
Assets
Cash 16400
Accounts receivable 1200
Office supplies 900
Equipment 1800
Furniture 4200
Total assets 24500
Liabilities and Stockholders' Equity
Accounts payable 4700
Unearned revenue 1400
Common stock 18000
Retained earnings 400
Total liabilities and stockholders' equity 24500

5. Return on assets = Net income/Average assets = $1800/$12250 = 14.69%

Average total assets = ($0 + $24500)/2 = $12250

Kindly round off the ROA as required.